# Kinematics of Circumgalactic OĀ VIĀ Gas and Disk Rotation of Star-forming Galaxies

### v1

@@ 2.3 Galaxy Properties
Note. ā (1) Name of the quasar. (2) Name of the galaxy. (3) Galaxy systemic redshift. (4) Galaxy stellar mass, rescaled to Chabrier IMF when necessary following Madau & Dickinson (2014). (5) Halo mass, derived using the stellar massāhalo mass relation in Behroozi etĀ al. (2010) except for those adopted from Kacprzak etĀ al. (2019). (6) Halo virial radius, defined by the overdensity in Bryan & Norman (1998) with respect to the critical density evaluated at the galaxy systemic redshift. (7) Asymptotic rotation speed of the galaxy disk. (8) Turnover radius from the arctangent rotation curve model. (9) Inclination angle of galactic disk. (10) Azimuthal angle, the angle between the quasar sightline and the galaxy major axis. (11) Sightline impact parameter, the projected distance between the centers of the galaxy and quasar. (12) Sightline impact parameter normalized by the halo virial radius.
@@ 2.3.2 Stellar Masses, Halo Masses, and Virial Radii
The stellar masses Ā of individual galaxies were obtained from previous work and were rescaled to the Chabrier initial mass function (IMF; Chabrier, 2003) following Madau & Dickinson (2014). Stellar masses extracted from Martin etĀ al. (2019) were in Chabrier IMF, whereas Werk etĀ al. (2012) assumed the Salpeter IMF (Salpeter, 1955), and Heckman etĀ al. (2017) adopted the Kroupa IMF (Kroupa, 2001) as in the MPA-JHU Value-Added Galaxy Catalog. Then, the halo mass Ā of each galaxy was calculated using the stellar massāhalo mass (SM-HM) relation derived from abundance matching in Behroozi etĀ al. (2010). The virial radius Ā was obtained following Ho etĀ al. (2017) using the overdensity defined by Bryan & Norman (1998).
@@
TableĀ 2 lists the stellar mass, halo mass, and virial radius of individual galaxies, and we label the references of the masses. The median Ā and Ā are 10.14 and 11.72, respectively. The Table also shows the sightline impact parameter normalized by the halo virial radius . All except one sightline intersect the galaxies within the virial radius (), and most of the sightlines lie within 0.5Ā from the galaxies (also see FigureĀ 1).
@@ 2.3.3 Galaxy Environments and Potential Interactions
However, the only exception is the target galaxy J091954+291345 at . We found that two red galaxies identified by the SHELS galaxy redshift survey (Geller etĀ al., 2014) have spectroscopic redshifts similar to our target. At the northeast of our target galaxy, a bright red galaxy 44 away (17 kpc) has a redshift of 0.23208, i.e., km s-1Ā relative to our galaxy. The other fainter, red galaxy has a redshift of 0.23179, i.e., km s-1, at 87 (33 kpc) south of our target. Hence, these three galaxies form a group with velocity differences within a few hundred km s-1. While we present the spectroscopic measurements for this target galaxy and the absorption lines measured in the quasar sightline, we exclude this system from the CGM kinematics analyses in later sections.
@@ 2.4 Quasar Spectroscopy with HST/COS
For each quasar, the raw data from COS were processed by the CALCOS pipeline (v3.3.9), and the one-dimensional spectra (x1d files) were retrieved from MAST444http://archive.stsci.edu. Custom software was used to align the one-dimensional spectra using the Milky Way absorption lines and coadd the aligned spectra. Because COS FUV spectra were oversampled, the coadded spectra were binned by three pixels to increase the signal-to-noise ratio. Finally, each coadded spectrum was continuum-normalized by fitting the quasar continuum with an Akime Spline at the absorption-free regions using the linetools555https://github.com/linetools/linetools python package.
@@ 2.5 Absorption-line Measurements: Voigt Profile Fitting and Column-density-weighted Velocity
We identified absorption-line systems within km s-1Ā of the systemic velocity of the target galaxies. We modeled individual velocity components with Voigt profiles and adopted the atomic data from Morton (2003). The modeled Voigt profile characterized the velocity centroid , the line width, i.e., the Doppler parameter , and the column density of individual velocity component. Using the VoigtFit python package (Krogager, 2018), we convolved the models with the COS line-spread function corresponding to the COS lifetime position when the observation was conducted. We compared the convolved line profiles to the data, and different transitions of the same ionic species were fitted jointly. For each ionic species, we created different sets of models with different number of velocity components, and the best-fit solution of each model set was determined using a minimization algorithm. Finally, we adopted the best-fit model that required the minimum number of velocity components to produce a reasonable fit and with the reduced- value closed to one.
@@
The right column of FigureĀ 2 shows the OĀ VIĀ absorption lines relative to the systemic redshifts of individual target galaxies. The fitted Voigt profiles of individual velocity components and the overall OĀ VIĀ absorption are shown in blue dashed and solid curves, respectively, and the blue ticks (top) mark the velocity centroids of individual components. TableĀ 4 lists the parameters of the Voigt profiles of the fitted OĀ VIĀ absorption.666 The Figures of the fitted line profiles of other ions and Voigt profile parameters are shown in FigureĀ 13 and TableĀ 6, respectively, in the Appendix. Only the J103640565125 sightline did not detect OĀ VIĀ around the target galaxy. In that case, we list the column density upper limit; this was derived from the linear section of the curve-of-growth and the upper limit of the rest-frame equivalent width of OĀ VIĀ 1031, measured using a velocity window that was three times the resolution element centered at the galaxy systemic velocity.
@@
Excluded from OĀ VIĀ kinematics analysis due to group environment | ||||||
@@
adding a random value within their 1- error bars,
@@
where indicates the -th velocity component, and and represent the perturbed velocity centroid and column density, respectively. From the resultant distribution of the weighted velocity, we adopt the median as the column-density-weighted velocity and the 68% confidence interval as the uncertainties. We show the OĀ VIĀ column-density-weighted velocity Ā for each sightline in the rightmost column of TableĀ 4.
@@ 3 CGM Kinematics and Galaxy Disk Rotation
With our sample of quasar sightlines near the major axes of galaxies, we examine the relationship between CGM kinematics and galaxy disk rotation. SectionĀ 3.1 and SectionĀ 3.2 focus on the OĀ VIĀ CGM and the low-ionization-state gas, respectively. In SectionĀ 3.3, we then compare the kinematics of the low ions and OĀ VIĀ detected in individual sightlines. In our analyses, we exclude galaxy J091954+291345 due to its close group environment (SectionĀ 2.3.3) and focus on the remaining 17 galaxyāquasar pairs.
3.1 Kinematics of OĀ VIĀ Gas
Among the 17 galaxyāquasar pairs, all sightlines except J103640+565125 have detected OĀ VIĀ around the host galaxies. FigureĀ 5 compares the kinematics of the OĀ VIĀ gas and that of the galaxy disk rotation. The circles and the dashed curves represent the velocity centroids of individual OĀ VIĀ velocity components and the galaxy rotation curves, respectively. With similar numbers of OĀ VIĀ velocity components that corotate (21) and counter-rotate (15) with the disk (TableĀ 5), it seems to suggest a lack of correlation between disk rotation and the kinematics of individual OĀ VIĀ velocity components. However, except for the sightline at the highest impact parameter of 284 kpc, the column-density-weighted velocities (black crosses) suggest that the bulk of OĀ VIĀ gas either corotates with the disk or is consistent with being at the systemic velocity of the host galaxies (within the grey band).777 Velocities smaller than the measurement uncertainties of the galaxy redshift of 25 km s-1Ā are considered as being consistent with the galaxy systemic velocity. In other words, we rarely found sightlines with net counter-rotating OĀ VIĀ absorbers.
In the bottom panel of FigureĀ 5, we normalized the velocities and impact parameters by the asymptotic rotation speeds and virial radii of the galaxies, respectively. The velocity centroids of most OĀ VIĀ components (34 out of 41) are smaller than the asymptotic rotation speed, i.e., . Sightlines with OĀ VIĀ components exceeding Ā typically have large impact parameters of 0.5. The only exception is the sightline around host galaxy J160951+353838, but this galaxy has a tail-like structure and an abnormally low asymptotic rotation speed of 45 km s-1Ā (SectionĀ 2.3.3). This explains the high values even though the measured OĀ VIĀ velocities in km s-1Ā are comparable to those in other sightlines.
OĀ VIĀ Category | Kinematics Class | Ratio | |
@@ 3.1 Kinematics of OĀ VIĀ Gas
Sightlines with SiĀ IIĀ and SiĀ IIIĀ Spectral Coverage ()ā | |||
@@
Counter-rotate | 8 | 0.53 |
Note. ā (1) Category of OĀ VIĀ velocity components. (2) Classification of OĀ VIĀ velocity components based on their Doppler sign relative to the disk rotation at the quasar side of the galaxy disk. (3) Number of OĀ VIĀ velocity components. (4) The ratio of the number of OĀ VIĀ velocity components between the kinematics class and the total within the same OĀ VIĀ category.
@@ 3.2 Corotation of the Low-ionization-state Gas with the Disk
Low-ionization-state ions trace cooler gas compared to OĀ VIĀ due to their lower ionization potentials (IP), e.g., SiĀ IIĀ and SiĀ IIIĀ with IP of 8.2 eV and 16.3 eV, respectively, compared to 113.9 eV for producing OĀ VI. For our sample, only 13 sightlines have COS spectra that cover the redshifted wavelengths of SiĀ II888 We focus on the SiĀ II 1190, 1193Ā and SiĀ II 1260Ā transitions, because other SiĀ IIĀ transition lines within the spectral coverage are either too weak to be detected or blended with other absorption lines. and SiĀ IIIĀ of the host galaxies, including the J103640+565125 sightline without OĀ VIĀ detection. Among these 13 sightlines, SiĀ IIIĀ exhibits a higher detection rate compared to SiĀ II, with 7 and 11 sightlines showing SiĀ IIĀ and SiĀ IIIĀ detection, respectively.
@@
FigureĀ 6 compares the kinematics of the SiĀ IIĀ (left) and SiĀ IIIĀ (right) gas to disk rotation. Except for the sightline at 184 kpc from galaxy J084356+261855, six out of the seven SiĀ II-detected sightlines show net corotating SiĀ IIĀ absorption, and only two (out of 21) SiĀ IIĀ velocity components counter-rotate with the disk. Net counter-rotating SiĀ IIIĀ absorbers are also rare, but the fraction of SiĀ IIIĀ velocity components that counter-rotate with the disk becomes higher compared to that of SiĀ II.
@@
FigureĀ 7 shows the fraction of non-counter-rotating velocity components of SiĀ II, SiĀ III, and OĀ VI, i.e., the gas either corotates with the disk or consistent with being at the systemic velocity of the host galaxy.999 Both black hatched and maroon filled histograms represent OĀ VI, but the former includes all sightlines, whereas the latter only includes the 13 sightlines with SiĀ IIĀ and SiĀ IIIĀ spectral coverage. Considering all sightlines regardless of their impact parameters (āAllā), SiĀ IIĀ has the highest fraction of non-counter-rotating velocity components of 90%. The fraction drops with increasing ionization potential, for which only 75% (60%) of the SiĀ IIIĀ (OĀ VI) velocity components are not counter-rotating. If we only focus on the sightlines with impact parameters below 100 kpc, the fraction of non-counter-rotating velocity components for all three ions become comparable (between 75% and 95%). However, in sightlines beyond 100 kpc from the host galaxies, OĀ VIĀ shows a significantly lower non-counter-rotating fraction compared to SiĀ IIĀ and SiĀ IIIĀ in the same impact parameter range and also compared to the OĀ VIĀ detected at small impact parameters. Therefore, in contrast to low-ionization-state gas that rarely counter-rotates with the disk, whether individual velocity components of the OĀ VIĀ gas counter-rotate depends on the sightline impact parameter, with sightlines at larger impact parameters more frequently detect counter-rotating OĀ VIĀ gas.
3.3 Comparison Between OĀ VIĀ and Low-ionization-state Gas Kinematics
@@
SectionsĀ 3.1 and 3.2 have shown that the low-ionization-state gas traced by SiĀ IIĀ and SiĀ IIIĀ rarely counter-rotate with the disk compared to OĀ VI, suggesting that OĀ VIĀ gas exhibits different kinematics compared to the lower ionization counterparts. In FigureĀ 8, we compare the Voigt profile fits of OĀ VIĀ to those of SiĀ IIĀ and SiĀ IIIĀ measured in individual sightlines.101010 For the J154741343357 sightline, the SiĀ IIIĀ absorption-line profile is incomplete, because it falls at the reddest end of the spectrum. The J103640+565125 sightline is not shown due to OĀ VIĀ non-detection. We arrange the panels in increasing impact parameter of the quasar sightlines. First, SiĀ IIĀ and SiĀ IIIĀ are more frequently detected in sightlines at low impact parameters, and the two sightlines beyond 200 kpc from the host galaxies neither detect SiĀ IIĀ nor SiĀ III. Secondly, the line profiles of different ions have complex component structures, especially sightlines closer to the galaxies tend to have more velocity components. While some OĀ VIĀ components have velocities comparable to their low-ionization-state counterparts, some do not affiliate with the low ions. But regardless, the overall absorption of different ions typically occupy similar velocity ranges.
@@
Motivated by FigureĀ 8 that shows that only some OĀ VIĀ components have velocities comparable to those of the low ions, we examine how the kinematics association between OĀ VIĀ and low ions (or the lackthereof) relates to the comparison of OĀ VIĀ kinematics and disk rotation. In each sightline, we attempt to match individual OĀ VIĀ velocity components with those of SiĀ IIĀ and SiĀ IIIĀ by minimizing the difference of their velocity centroids. We consider an OĀ VIĀ velocity component to be successfully matched with a SiĀ IIĀ or SiĀ IIIĀ component if the velocity difference is under 30 km s-1, a threshold that includes the uncertainties in defining the galaxy systemic velocity, the wavelength calibration of the COS spectra111111 According to the āCycle 26 COS/FUV Spectroscopic Sensitivity Monitor (ISR 2020-06)ā, the COS wavelength zeropoint uncertainties are 3 pixels, i.e., km s-1. , and the typical uncertainty of the velocity centroid from Voigt profile fitting.121212 While increasing (decreasing) the 30 km s-1Ā threshold value will increase (reduce) the number of low-ion matched OĀ VIĀ velocity components, it does not affect the main conclusions from the kinematics comparison between OĀ VIĀ components with and without low-ion matches. Each SiĀ IIĀ or SiĀ IIIĀ velocity component can only be matched with one OĀ VIĀ component (and vice versa). Hence, if more than one SiĀ IIĀ or SiĀ IIIĀ component satisfy the matching criteria with OĀ VI, then we select the one with the smallest velocity difference.
@@
Analogous to FigureĀ 5, FigureĀ 9 compares the OĀ VIĀ kinematics with galaxy disk rotation by showing the 12 sightlines with OĀ VIĀ detection and spectral coverage of SiĀ IIĀ and SiĀ III. Each marker represents the velocity centroid of an OĀ VIĀ component, and the yellow diamonds (top) and cyan squares (bottom) highlight the SiĀ II- and SiĀ III-matched OĀ VIĀ components, respectively. Both panels clearly show that SiĀ II- or SiĀ III-matched OĀ VIĀ velocity components rarely counter-rotate with the galaxy disk (also see TableĀ 5). In fact, none of the SiĀ II-matched OĀ VIĀ velocity components counter-rotates (0/10), in contrast to the 10/19 (53%) no-SiĀ II-matched OĀ VIĀ components that counter-rotate with the disk. While counter-rotating SiĀ III-matched OĀ VIĀ components exist (2/14; 14%), that occurs less frequently compared to the OĀ VIĀ components without a SiĀ IIIĀ matched (8/15; 53%).

### v2

@@ 2.3 Galaxy Properties
Note. ā (1) Name of the quasar. (2) Name of the galaxy. (3) Galaxy systemic redshift. The typical uncertainty is 25 km s-1. (4) Galaxy stellar mass, rescaled to Chabrier IMF when necessary following Madau & Dickinson (2014). The typical uncertainty is 0.2 dex. (5) Halo mass, derived using the stellar massāhalo mass relation in Behroozi etĀ al. (2010) except for those adopted from Kacprzak etĀ al. (2019). The typical uncertainty of halo mass is 0.2 dex; this does not include the intrinsic scatter in stellar mass at a single halo mass from the mean stellar massāhalo mass relation. (6) Halo virial radius, defined by the overdensity in Bryan & Norman (1998) with respect to the critical density evaluated at the galaxy systemic redshift. The virial radius has a typical uncertainty of 15 kpc. (7) Asymptotic rotation speed of the galaxy disk. The typical uncertainty is 15 km s-1. (8) Turnover radius from the arctangent rotation curve model. The turnover radius has a typical uncertainty of 0.3 kpc. (9) Inclination angle of galactic disk. The typical uncertainty is 3ā. (10) Azimuthal angle, the angle between the quasar sightline and the galaxy major axis. The typical uncertainty of the azimuthal angle is 2ā, which is dominated by the uncertainty of the position angle of the galaxy major axis. (11) Sightline impact parameter, the projected distance between the centers of the galaxy and quasar. The typical uncertainty of impact parameter is less than 1 kpc, which comes from the uncertainty in the angular separation between the galaxy and the quasar. (12) Sightline impact parameter normalized by the halo virial radius. The typical uncertainty is 0.07, dominated by the uncertainty in the virial radius.
@@ 2.3.2 Stellar Masses, Halo Masses, and Virial Radii
The stellar masses of individual galaxies were obtained from previous work and were rescaled to the Chabrier initial mass function (IMF; Chabrier, 2003) following Madau & Dickinson (2014). Stellar masses extracted from Martin etĀ al. (2019) were in Chabrier IMF, whereas Werk etĀ al. (2012) assumed the Salpeter IMF (Salpeter, 1955), and Heckman etĀ al. (2017) adopted the Kroupa IMF (Kroupa, 2001) as in the MPA-JHU Value-Added Galaxy Catalog. Then, the halo mass of each galaxy was calculated using the stellar massāhalo mass (SM-HM) relation derived from abundance matching in Behroozi etĀ al. (2010). The virial radius was obtained following Ho etĀ al. (2017) using the overdensity defined by Bryan & Norman (1998).
@@
TableĀ 2 lists the stellar mass, halo mass, and virial radius of individual galaxies, and we label the references of the masses. The median and are 10.14 and 11.72, respectively. The Table also shows the sightline impact parameter normalized by the halo virial radius . All except one sightline intersect the galaxies within the virial radius (), and most of the sightlines lie within 0.5 from the galaxies (also see FigureĀ 1).
@@ 2.3.3 Galaxy Environments and Potential Interactions
However, the only exception is the target galaxy J091954+291345 at . We found that two red galaxies identified by the SHELS galaxy redshift survey (Geller etĀ al., 2014) have spectroscopic redshifts similar to our target. At the northeast of our target galaxy, a bright red galaxy 44 away (17 kpc) has a redshift of 0.23208, i.e., km s-1 relative to our galaxy. The other fainter, red galaxy has a redshift of 0.23179, i.e., km s-1, at 87 (33 kpc) south of our target. Hence, these three galaxies form a group with velocity differences within a few hundred km s-1. While we present the spectroscopic measurements for this target galaxy and the absorption lines measured in the quasar sightline, we exclude this system from the CGM kinematics analyses in later sections.
@@ 2.4 Quasar Spectroscopy with HST/COS
For each quasar, the raw data from COS were processed by the CALCOS pipeline (v3.3.9), and the one-dimensional spectra (x1d files) were retrieved from MAST444http://archive.stsci.edu. Custom software was used to align the one-dimensional spectra using the Milky Way absorption lines and coadd the aligned spectra. Because COS FUV spectra were oversampled, the coadded spectra were binned by three pixels to increase the signal-to-noise ratio. Finally, each coadded spectrum was continuum-normalized by fitting the quasar continuum with an Akime Spline at the absorption-free regions using the linetools555https://github.com/linetools/linetools python package (Prochaska etĀ al., 2017).
@@ 2.5 Absorption-line Measurements: Voigt Profile Fitting and Column-density-weighted Velocity
We identified absorption-line systems within km s-1 of the systemic velocity of the target galaxies. We modeled individual velocity components with Voigt profiles and adopted the atomic data from Morton (2003). The modeled Voigt profile characterized the velocity centroid , the line width, i.e., the Doppler parameter , and the column density of individual velocity component. Using the VoigtFit python package (Krogager, 2018), we convolved the models with the COS line-spread function corresponding to the COS lifetime position when the observation was conducted. We compared the convolved line profiles to the data, and different transitions of the same ionic species were fitted jointly. For each ionic species, we created different sets of models with different number of velocity components, and the best-fit solution of each model set was determined using a minimization algorithm. Finally, we adopted the best-fit model that required the minimum number of velocity components to produce a reasonable fit and with the reduced- value closed to one.
@@
The right column of FigureĀ 2 shows the OĀ VI absorption lines relative to the systemic redshifts of individual target galaxies. The fitted Voigt profiles of individual velocity components and the overall OĀ VI absorption are shown in blue dashed and solid curves, respectively, and the blue ticks (top) mark the velocity centroids of individual components. TableĀ 4 lists the parameters of the Voigt profiles of the fitted OĀ VI absorption.666The Figures of the fitted line profiles of other ions and Voigt profile parameters are shown in FigureĀ 13 and TableĀ 6, respectively, in the Appendix. Only the J103640565125 sightline did not detect OĀ VI around the target galaxy. In that case, we list the column density upper limit; this was derived from the linear section of the curve-of-growth and the upper limit of the rest-frame equivalent width of OĀ VI 1031, measured using a velocity window that was three times the resolution element centered at the galaxy systemic velocity.
It is worth noting that in three sightlines, both OĀ VI 1031, 1037 transitions include intervening absorption at the same velocity range relative to the systemic velocity of our target galaxies. At +140 km s-1 from the galaxy in the J113910135043 sightline, the OĀ VI 1031 absorption is overlapped with the Ly absorption at , whereas OĀ VI 1037 is blended with SĀ II of the Milky Way (MW) absorption, Ly at , and Ly at . We masked OĀ VI 1037 due to the complexity of absorption from multiple species at different redshifts. As for OĀ VI 1031, we first attempted to perform Voigt profile fitting together with the HĀ I absorption at , but that resulted in huge uncertainties in the Voigt profile parameters. So instead, we fitted the HĀ I lines at alone (from Ly to Ly), through which we determined that the Ly absorption fully accounted for the +140 km s-1 absorption near the OĀ VI 1031 line of our target galaxy. We then fixed the Voigt profile parameters of the Ly absorption line while fitting the OĀ VI 1031, 1037 doublet for our target galaxy. For the J132222+464546 sightline, the absorption at +100 km s-1 of OĀ VI 1031, 1037 is blended with the SĀ II 1253 and SiĀ II 1260 absorption of the MW, respectively. We masked the absorption of OĀ VI 1037 and MW SiĀ II 1260 due to the saturation of the MW SiĀ II absorption lines. Then, we fitted the OĀ VI 1031 and the MW SĀ II 1253, 1259 absorption together, for which the MW SĀ II 1253 fully accounted for the extra absorption at +100 km s-1 from our galaxy. As for the J160951+353843 sightline, while the OĀ VI doublet showed absorption at 250 km s-1, that part of the absorption in the OĀ VI 1037 panel was identified as the CĀ II 1036 absorption of our target galaxy. Nevertheless, for all three cases, if part of the intervening absorption actually includes OĀ VI contribution from our target galaxies, then the bulk of OĀ VI absorption would become more Doppler shifted towards the same direction as the disk rotation, which thereby strengthened our conclusion that the bulk of OĀ VI gas rarely counter-rotates (SectionĀ 3).
@@
| Excluded from OĀ VI kinematics analysis due to group environment | ||||||
@@
adding a random value drawn from
a Gaussian distribution with standard deviation
equals to
their 1- error bars,
@@
where indicates the -th velocity component, and and represent the perturbed velocity centroid and column density, respectively. From the resultant distribution of the weighted velocity, we adopt the median as the column-density-weighted velocity and the 68% confidence interval as the uncertainties. We show the OĀ VI column-density-weighted velocity for each sightline in the rightmost column of TableĀ 4.
@@ 3 CGM Kinematics and Galaxy Disk Rotation
With our sample of quasar sightlines near the major axes of galaxies, we examine the relationship between CGM kinematics and galaxy disk rotation. SectionĀ 3.1 and SectionĀ 3.2 focus on the OĀ VI CGM and the low-ionization-state gas, respectively. In SectionĀ 3.3, we then compare the kinematics of the low ions and OĀ VI detected in individual sightlines. In our analyses, we exclude galaxy J091954+291345 due to its close group environment (SectionĀ 2.3.3) and focus on the remaining 17 galaxyāquasar pairs.
3.1 Kinematics of OĀ VI Gas
Among the 17 galaxyāquasar pairs, all sightlines except J103640+565125 have detected OĀ VI around the host galaxies. FigureĀ 5 compares the kinematics of the OĀ VI gas and that of the galaxy disk rotation. The circles and the dashed curves represent the velocity centroids of individual OĀ VI velocity components and the galaxy rotation curves, respectively. With similar numbers of OĀ VI velocity components that corotate (21) and counter-rotate (15) with the disk (TableĀ 5), it seems to suggest a lack of correlation between disk rotation and the kinematics of individual OĀ VI velocity components. However, except for the sightline at the highest impact parameter of 284 kpc, the column-density-weighted velocities (black crosses) suggest that the bulk of OĀ VI gas either corotates with the disk or is consistent with being at the systemic velocity of the host galaxies (within the grey band).777Velocities smaller than the measurement uncertainties of the galaxy redshift of 25 km s-1 are considered as being consistent with the galaxy systemic velocity. In other words, we rarely found sightlines with net counter-rotating OĀ VI absorbers.
In the bottom panel of FigureĀ 5, we normalized the velocities and impact parameters by the asymptotic rotation speeds and virial radii of the galaxies, respectively. The velocity centroids of most OĀ VI components (34 out of 41) are smaller than the asymptotic rotation speed, i.e., . Sightlines with OĀ VI components exceeding typically have large impact parameters of 0.5. The only exception is the sightline around host galaxy J160951+353838, but this galaxy has a tail-like structure and an abnormally low asymptotic rotation speed of 45 km s-1 (SectionĀ 2.3.3). This explains the high values even though the measured OĀ VI velocities in km s-1 are comparable to those in other sightlines.
| OĀ VI Category | Kinematics Class | Ratio | |
@@ 3.1 Kinematics of OĀ VI Gas
| Sightlines with SiĀ II and SiĀ III Spectral Coverage ()ā | |||
@@
| Counter-rotate | 8 | 0.53 | |
Note. ā (1) Category of OĀ VI velocity components. (2) Classification of OĀ VI velocity components based on their Doppler sign relative to the disk rotation at the quasar side of the galaxy disk. (3) Number of OĀ VI velocity components. (4) The ratio of the number of OĀ VI velocity components between the kinematics class and the total within the same OĀ VI category.
@@ 3.2 Corotation of the Low-ionization-state Gas with the Disk
Low-ionization-state ions trace cooler gas compared to OĀ VI due to their lower ionization potentials (IP), e.g., SiĀ II and SiĀ III with IP of 8.2 eV and 16.3 eV, respectively, compared to 113.9 eV for producing OĀ VI. For our sample, only 13 sightlines have COS spectra that cover the redshifted wavelengths of SiĀ II888We focus on the SiĀ II 1190, 1193 and SiĀ II 1260 transitions, because other SiĀ II transition lines within the spectral coverage are either too weak to be detected or blended with other absorption lines. and SiĀ III of the host galaxies, including the J103640+565125 sightline without OĀ VI detection. Among these 13 sightlines, SiĀ III exhibits a higher detection rate compared to SiĀ II, with 7 and 11 sightlines showing SiĀ II and SiĀ III detection, respectively.
@@
FigureĀ 6 compares the kinematics of the SiĀ II (left) and SiĀ III (right) gas to disk rotation. Except for the sightline at 184 kpc from galaxy J084356+261855, six out of the seven SiĀ II-detected sightlines show net corotating SiĀ II absorption, and only two (out of 21) SiĀ II velocity components counter-rotate with the disk. Net counter-rotating SiĀ III absorbers are also rare, but the fraction of SiĀ III velocity components that counter-rotate with the disk becomes higher compared to that of SiĀ II.
3.3 Comparison Between OĀ VI and Low-ionization-state Gas Kinematics
@@
SectionsĀ 3.1 and 3.2 have shown that the low-ionization-state gas traced by SiĀ II and SiĀ III rarely counter-rotate with the disk compared to OĀ VI, suggesting that OĀ VI gas exhibits different kinematics compared to the lower ionization counterparts. In FigureĀ 7, we show the fraction of non-counter-rotating velocity components of SiĀ II, SiĀ III, and OĀ VI, i.e., the gas either corotates with the disk or consistent with being at the systemic velocity of the host galaxy.999Both black hatched and maroon filled histograms represent OĀ VI, but the former includes all sightlines, whereas the latter only includes the 13 sightlines with SiĀ II and SiĀ III spectral coverage. The error bars represent the Poisson confidence limits calculated from the number of non-counter-rotating velocity components, and then divided by the total number of velocity components of the same ion. Considering all sightlines regardless of their impact parameters (āAllā), SiĀ II has the highest fraction of non-counter-rotating velocity components of 90%. The fraction drops with increasing ionization potential, for which only 75% (60%) of the SiĀ III (OĀ VI) velocity components are not counter-rotating. If we only focus on the sightlines with impact parameters below 100 kpc, the fraction of non-counter-rotating velocity components for all three ions become comparable (between 75% and 95%). However, in sightlines beyond 100 kpc from the host galaxies, OĀ VI shows a lower non-counter-rotating fraction compared to SiĀ II and SiĀ III in the same impact parameter range and also compared to the OĀ VI detected at small impact parameters. Therefore, in contrast to low-ionization-state gas that rarely counter-rotates with the disk, whether individual velocity components of the OĀ VI gas counter-rotate possibly depends on the sightline impact parameter, with sightlines at larger impact parameters more frequently detect counter-rotating OĀ VI gas.
@@
FigureĀ 8 compares the Voigt profile fits of OĀ VI to those of SiĀ II and SiĀ III measured in individual sightlines.101010For the J154741343357 sightline, the SiĀ III absorption-line profile is incomplete, because it falls at the reddest end of the spectrum. The J103640+565125 sightline is not shown due to OĀ VI non-detection. We arrange the panels in increasing impact parameter of the quasar sightlines. First, SiĀ II and SiĀ III are more frequently detected in sightlines at low impact parameters, and the two sightlines beyond 200 kpc from the host galaxies neither detect SiĀ II nor SiĀ III. Secondly, the line profiles of different ions have complex component structures, especially sightlines closer to the galaxies tend to have more velocity components. While some OĀ VI components have velocities comparable to their low-ionization-state counterparts, some do not affiliate with the low ions. But regardless, the overall absorption of different ions typically occupy similar velocity ranges.
@@
Motivated by FigureĀ 8 that shows that only some OĀ VI components have velocities comparable to those of the low ions, we examine how the kinematics association between OĀ VI and low ions (or the lackthereof) relates to the comparison of OĀ VI kinematics and disk rotation. In each sightline, we attempt to match individual OĀ VI velocity components with those of SiĀ II and SiĀ III by minimizing the difference of their velocity centroids. We consider an OĀ VI velocity component to be successfully matched with a SiĀ II or SiĀ III component if the velocity difference is under 30 km s-1, a threshold that includes the uncertainties in defining the galaxy systemic velocity, the wavelength calibration of the COS spectra111111According to the āCycle 26 COS/FUV Spectroscopic Sensitivity Monitor (ISR 2020-06)ā, the COS wavelength zeropoint uncertainties are 3 pixels, i.e., km s-1. , and the typical uncertainty of the velocity centroid from Voigt profile fitting.121212While increasing (decreasing) the 30 km s-1 threshold value will increase (reduce) the number of low-ion matched OĀ VI velocity components, it does not affect the main conclusions from the kinematics comparison between OĀ VI components with and without low-ion matches. Each SiĀ II or SiĀ III velocity component can only be matched with one OĀ VI component (and vice versa). Hence, if more than one SiĀ II or SiĀ III component satisfy the matching criteria with OĀ VI, then we select the one with the smallest velocity difference.
@@
Analogous to FigureĀ 5, FigureĀ 9 compares the OĀ VI kinematics with galaxy disk rotation by showing the 12 sightlines with OĀ VI detection and spectral coverage of SiĀ II and SiĀ III. Each marker represents the velocity centroid of an OĀ VI component, and the yellow diamonds (top) and cyan squares (bottom) highlight the SiĀ II- and SiĀ III-matched OĀ VI components, respectively. Both panels clearly show that SiĀ II- or SiĀ III-matched OĀ VI velocity components rarely counter-rotate with the galaxy disk (also see TableĀ 5). In fact, none of the SiĀ II-matched OĀ VI velocity components counter-rotates (0/10), in contrast to the 10/19 (53%) no-SiĀ II-matched OĀ VI components that counter-rotate with the disk. While counter-rotating SiĀ III-matched OĀ VI components exist (2/14; 14%), that occurs less frequently compared to the OĀ VI components without a SiĀ III matched (8/15; 53%). The fact that half of the unmatched OĀ VI components counter-rotate also implies a true distinction in the OĀ VI kinematics between velocity components with and without low-ion matches. Consider a hypothetical scenario that OĀ VI is randomly moving and has no correlation with the kinematics of the low ions at all. Then we should find corotating and counter-rotating OĀ VI components each at 50% of the time. When we match OĀ VI components with SiĀ II and SiĀ III by minimizing their velocity differences, we will find several OĀ VI components aligning with the low ions by chance. Because low ions tend to corotate with the disk (SectionĀ 3.2), the OĀ VI components with low-ion matches will also tend to corotate, whereas the remaining unmatched OĀ VI components will generally counter-rotate. However, our measurements show that only about 50% of the unmatched OĀ VI components counter-rotate. Therefore, this invalidates the hypothetical scenario, indicating that the successful matches between OĀ VI and low ions did not occur by chance. In other words, while the low-ion matched OĀ VI components rarely counter-rotate, the unmatched OĀ VI components do not preferentially corotate nor counter-rotate with the disk.